OpenStudy (anonymous):

Which is the equation of the line that contains points (0, 5) and (5, 8)?

OpenStudy (luigi0210):

Start by finding the slope then use the point-slope formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(8 - 5)/(5 - 0) = 3/5

OpenStudy (geneticrockhopper247):

I think that there are supposed to be an x and y in point slope form. Let me brush up on it right quick and then I can check this.

OpenStudy (geneticrockhopper247):

The formula for this is y-y1=m(x-x1). Do you know what this means @sandrayanez ?